# Dynamic and Thermodynamic Stability of Superconducting-superfluid Stars

### v1

@@ V.3 Trivial displacements
so $\bm{H}^{\text{X}}$ may be viewed as a 2-form on the manifold of orbits of $u_{\text{X}}^{a}$. Assuming that the our star is simply connected, $d\bm{H}^{\text{X}}=0$ implies that
@@ VI Canonical energy and dynamic stability
It is easy to show that not only is the canonical energy symmetric and conserved, but also gauge invariant in the sense that
$\mathcal{E}\left(\delta_{1}\phi,\delta_{2}\phi\right)=\mathcal{E}\left(\delta_{1}\phi,\delta_{2}\phi+\mathscr{L}_{X}\phi\right),$ (193)
with $X^{a}$ smooth and of compact support. Moreover, since our star has a spatial compact support, then $\mathcal{E}\left(\delta\phi,\delta\phi\right)$ has a non-negative net flux at null infinity if the perturbation is asymptotically physically stationary (which will be defined below) at late times. This has been shown in [Shi:2022jya].
A smooth linearized solution $\delta\phi=\left(\delta g_{ab},\delta A_{a},\xi_{\text{X}}^{a}\right)$ is said to be physically stationary if the physical fields $\delta g_{ab}$, $\delta A_{a}$, $\delta\bm{N}^{\text{X}}$, and $\delta s$ can be made stationary by a gauge transformation, i.e., if there exists a smooth vector field $X^{a}$, which is an asymptotic symmetry near infinity, such that
$0$ $=\mathscr{L}_{t}\left(\delta g_{ab}+\mathscr{L}_{X}g_{ab}\right),$ (194)
$0$ $=\mathscr{L}_{t}\left(\delta A_{a}+\mathscr{L}_{X}A_{a}\right),$ (195)
$0$ $=\mathscr{L}_{t}\left(\delta\bm{N}^{\text{X}}+\mathscr{L}_{X}\bm{N}^{\text{X}}\right)=-\mathscr{L}_{\left[t,\xi_{\text{X}}-X\right]}\bm{N}^{\text{X}},$ (196)
$0$ $=\mathscr{L}_{t}\left(\delta s+\mathscr{L}_{X}s\right)=-\mathscr{L}_{\left[t,\xi_{\text{X}}-X\right]}s.$ (197)
Note that the last two equations Eqs. (196) and (196) means that the perturbation $\left(0,0,\left[t,\xi_{\text{X}}-X\right]^{a}\right)$ is trivial perturbation with trivial displacements $\left[t,\xi_{\text{X}}-X\right]^{a}$. So equivalently, a linearized solution $\delta\phi$ is physically stationary if and only if there exists a smooth vector field $X^{a}$, which is an asymptotic symmetry near infinity, such that
$\mathscr{L}_{t}\delta\phi=$ $\left(-\mathscr{L}_{\left[t,X\right]}g_{ab},-\mathscr{L}_{\left[t,X\right]}A_{a},\zeta_{\text{X}}^{a}=\left[t,X\right]^{a}\right)$
@@
Unfortunately, we will find that $\mathcal{E}$ is not degenerate on all physically stationary solutions. In fact, since $\mathcal{E}\left(\delta^{\prime}\phi,\delta\phi\right)=W\left(\delta^{\prime}\phi,\mathscr{L}_{t}\delta\phi\right)$, it follows that $\delta\phi$ is a degeneracy of $\mathcal{E}$ if and only if $\mathscr{L}_{t}\delta\phi$ is a degeneracy of $W$. As discussed at the end of Sec. V.2, when restricted to $\mathcal{C}$, $W$ is degenerate precisely on the gauge transformations $\mathscr{L}_{X}\phi$ with $X^{a}$ smooth and going to zero at infinity , so that $\delta\phi$ in the domain of $\mathcal{E}$ is a degeneracy if and only if
$\mathscr{L}_{t}\delta\phi=\left(\mathscr{L}_{X}g_{ab},\mathscr{L}_{X}A_{a},\xi_{\text{X}}^{a}=-X^{a}\right).$ (200)
@@
$\mathcal{E}\left(\delta\phi,\delta_{\text{ps}}\phi\right)=$ $-W\left[\delta\phi,\mathscr{L}_{\left[t,X\right]}\phi\right]$
@@
For a general asymptotic symmetry $X^{a}$, the commutator $\left[t,X\right]^{a}$ is, at most, an asymptotic translation (as occurs when $X^{a}$ is an asymptotic boost). Therefore, in order the first term of right side vanishes, we need to restrict $\delta\phi$ such that $\delta H_{\left[t,X\right]}=0$, where $\delta H_{\left[t,X\right]}$ is the ADM linear momentum (see Eq. (45)). This in fact does not impose a physical restriction on the perturbations as we can achieve this by addition of the action of an infinitesimal Lorentz boost on the background solution. On the other hands, to make the second term vanishes, we need to restrict $\delta\phi$ such that
@@
for all trivial perturbations. As a consequence, let us define a subspace $\mathcal{V}\subset\mathcal{C}$ of the linearized solutions composed of perturbations that have $\delta H_{\left[t,X\right]}=0$ and are symplectically orthogonal to all trivial perturbations, in other words, $\mathcal{V}$ is the symplectic complement of the subspace $\mathcal{W}$ spanned by the perturbations of the form $\mathscr{L}_{t}\delta_{\text{ps}}\phi$. Note that the double symplectic complement of $\mathcal{W}$ is itself according to Eq. (165), thus the symplectic complement of $\mathcal{V}$ in itself is $\mathcal{V}\cap\mathcal{W}$, which implies that the degeneracy of the canonical energy $\mathcal{E}$ is given precisely by the physically stationary solutions when restricted onto $\mathcal{V}$.
@@
Let us consider the restriction condition given in Eq. (202). For non-axisymmetric perturbations, this restricted condition is in fact same as the condition to make Eq. (185) vanish. If the superfluid neutrons and superconducting protons in our star is irrotational, i.e., their vorticity vanish $\bm{w}^{\Upsilon}=0$, and so that their total momentum covector should be of the form
@@
As shown in [Shi:2022jya], if we further focus onto the background in which $\nabla_{a}s\neq 0$, then the condition Eq. (202) does not lead to a real physical restriction to $\mathcal{V}$ for non-axisymmetric perturbations, since it can be achieved in the suitable background solution by adding a trivial perturbation. However, if there are vortex in our superconducting-superfluid star and so that the irrotation is violated, then whether the condition Eq. (202) will impose a physical restriction to $\mathcal{V}$ is still unknown to our knowledge.
@@
So we still have the perturbed conservation of angular momentum, i.e., $\delta\left(d\bm{J}^{\text{X}}\right)=d\left(\delta\bm{J}^{\text{X}}\right)=0$. Replacing $\bm{N}^{\text{X}}$ by $\bm{J}^{\text{X}}$, then an identical calculation, besides the fifth equality in Eq. (212) replaced by
@@
where the second equality follows from the symmetry of $\mathcal{E}$, and the last equality follows from $\mathscr{L}_{t}\delta\phi$ satisfies $\Delta_{\text{X}}\bm{J}^{\text{X}}=0$. So $\mathscr{L}_{t}^{2}\delta\phi$ is symplectically orthogonal to all axisymmetric trivial perturbations. Furthermore, for a smooth vector field $X^{a}$ which is an asymptotic symmetry near infinity, since
$W\left(\mathscr{L}_{\left[t,X\right]}\phi,\mathscr{L}_{t}^{2}\delta\phi\right)=$ $\mathcal{E}\left(\mathscr{L}_{\left[t,X\right]}\phi,\mathscr{L}_{t}\delta\phi\right)$
$=$ $\mathcal{E}\left(\mathscr{L}_{t}\delta\phi,\mathscr{L}_{\left[t,X\right]}\phi\right)$
$=$ $W\left(\mathscr{L}_{t}\delta\phi,\mathscr{L}_{\left[t,\left[t,X\right]\right]}\phi\right)$
@@
where we have used the fact that $Y^{a}=\left[t,\left[t,X\right]\right]^{a}$ vanishes at infinity for any asymptotic symmetry generator $X^{a}$. Thus, we have shown that $\mathscr{L}_{t}^{2}\delta\phi\in\mathcal{V}$. ∎
@@ VII Thermodynamic stability
As shown in [Green:2013ica], there can be a non-axisymmetric perturbation which is made to have $\sum_{\text{X}}\Omega_{\text{X}}\delta^{2}J^{\text{X}}>\delta^{2}M$, whence $\mathcal{E}^{\prime}<0$, in other words, all rotating stars are thermodynamically unstable with respect to the non-axisymmetric perturbations. So we will solely consider the axisymmetric perturbations within the Lagrangian displacement framework such that $\delta J^{\text{X}}=0$ below, and we finally show that in the axisymmetric case, the canonical energy $\mathcal{E}\left(\delta\phi,\delta\phi\right)$ coincides with $\mathcal{E}^{\prime}$.
@@
$=$ $\int_{\Sigma}\left\{\delta\bm{\omega}^{\left(EM\right)}\left[\delta\phi,\left(\mathscr{L}_{t}g_{ab},\mathscr{L}_{t}A_{a}\right)\right]+\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right\}$
$+\int_{\Sigma}\left[\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)-\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right]$
$=$ $\int_{\Sigma}\delta\bm{\omega}\left(\delta\phi,\mathscr{L}_{t}\phi\right)$
$+\int_{\Sigma}\left[\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)-\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right]$
$=$ $\int_{\Sigma}\delta\left\{\iota_{t}\bm{E}\cdot\delta\phi+\delta\bm{C}_{t}+d\left[\delta\bm{Q}_{t}-\iota_{t}\bm{\theta}\left(\phi;\delta\phi\right)\right]\right\}$
$+\int_{\Sigma}\left[\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)-\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right]$
@@
$\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)$
$=$ $\sum_{\text{X}}\left(\mathscr{L}_{t}\xi_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}-\xi_{\text{X}}^{a}\mathscr{L}_{t}\delta\bm{P}_{a}^{\text{X}}-\left[\xi_{\text{X}},\mathscr{L}_{t}\xi_{\text{X}}\right]^{a}\bm{P}_{a}^{\text{X}}\right),$ (228)
@@
$=$ $\sum_{\text{X}}\left(t^{a}\delta^{2}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{t}\xi_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}-\left[\xi_{\text{X}},\mathscr{L}_{t}\xi_{\text{X}}\right]^{a}\bm{P}_{a}^{\text{X}}\right).$ (230)
@@
$=$ $\sum_{\text{X}}\left[|v_{\text{X}}|u_{\text{X}}^{a}\delta^{2}\bm{P}_{a}^{\text{X}}-\Omega_{\text{X}}\varphi^{a}\delta^{2}\bm{P}_{a}^{\text{X}}\right]$
$-\sum_{\text{X}}\left[2\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\delta\bm{P}_{a}^{\text{X}}-\mathscr{L}_{\xi_{\text{X}}}\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\bm{P}_{a}^{\text{X}}\right].$ (233)
@@
$=$ $-\delta\left(\frac{1}{2}u_{\text{X}}^{a}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-q_{\text{X}b}^{a}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\bm{P}_{a}^{\text{X}}$
$-\left(u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-2u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$
$=$ $-\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+\mathscr{L}_{\xi_{\text{X}}}\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}$
$+\left(u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-2u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$
$=$ $\mathscr{L}_{\xi_{\text{X}}}\left(\frac{1}{2}u_{\text{X}}^{a}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-q_{\text{X}b}^{a}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\bm{P}_{a}^{\text{X}}$
$+\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$
$=$ $\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}-\mathscr{L}_{\xi_{\text{X}}}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}$
$-\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$

### v2

@@ V.4 Trivial displacements
so $\bm{H}^{\text{X}}$ may be viewed as a 2-form on the manifold of orbits of $u_{\text{X}}^{a}$. Assuming that there are no holes in our star, then $d\bm{H}^{\text{X}}=0$ implies that
@@ VI Canonical energy and dynamic stability
It is easy to show that not only is the canonical energy symmetric and conserved. Indeed, the symmetry comes from that
$\mathcal{E}\left(\delta_{1}\phi,\delta_{2}\phi\right)-\mathcal{E}\left(\delta_{2}\phi,\delta_{1}\phi\right)$
$=$ $W\left(\delta_{1}\phi,\mathscr{L}_{t}\delta_{2}\phi\right)+W\left(\mathscr{L}_{t}\delta_{1}\phi,\delta_{2}\phi\right)$
$=$ $\int_{\Sigma}\left[\bm{\omega}\left(\delta_{1}\phi,\mathscr{L}_{t}\delta_{2}\phi\right)+\bm{\omega}\left(\mathscr{L}_{t}\delta_{1}\phi,\delta_{2}\phi\right)\right]$
$=$ $\int_{\Sigma}\mathscr{L}_{t}\bm{\omega}\left(\delta_{1}\phi,\delta_{2}\phi\right)$
$=$ $\int_{\Sigma}d\left[\iota_{t}\bm{\omega}\left(\delta_{1}\phi,\delta_{2}\phi\right)\right]$
$=$ $0.$ (209)
where we have used $W$ is antisymmetric in the first step and $\bm{\omega}$ is closed for on-shell linearized solution in the fourth step. The conservation of $\mathcal{E}$ also comes from the closeness of $\omega$. Furthermore, the canonical energy is gauge invariant in the sense that
$\mathcal{E}\left(\delta_{1}\phi,\delta_{2}\phi\right)=\mathcal{E}\left(\delta_{1}\phi,\delta_{2}\phi+\mathscr{L}_{\mathcal{X}}\phi\right),$ (210)
with $\mathcal{X}^{a}$ smooth and of compact support. Moreover, since our star has a spatial compact support, then $\mathcal{E}\left(\delta\phi,\delta\phi\right)$ has a non-negative net flux at null infinity if the perturbation is asymptotically physically stationary (which will be defined below) at late times. This has been shown in [32].
A smooth linearized solution $\delta\phi=\left(\delta g_{ab},\delta A_{a},\xi_{\text{X}}^{a}\right)$ is said to be physically stationary if the physical fields $\delta g_{ab}$, $\delta A_{a}$, $\delta\bm{N}^{\text{X}}$, and $\delta s$ can be made stationary by a gauge transformation, i.e., if there exists a smooth vector field $\mathcal{X}^{a}$, which is an asymptotic symmetry near infinity, such that
$0$ $=\mathscr{L}_{t}\left(\delta g_{ab}+\mathscr{L}_{\mathcal{X}}g_{ab}\right),$ (211)
$0$ $=\mathscr{L}_{t}\left(\delta A_{a}+\mathscr{L}_{\mathcal{X}}A_{a}\right),$ (212)
$0$ $=\mathscr{L}_{t}\left(\delta\bm{N}^{\text{X}}+\mathscr{L}_{\mathcal{X}}\bm{N}^{\text{X}}\right)=-\mathscr{L}_{\left[t,\xi_{\text{X}}-\mathcal{X}\right]}\bm{N}^{\text{X}},$ (213)
$0$ $=\mathscr{L}_{t}\left(\delta s+\mathscr{L}_{\mathcal{X}}s\right)=-\mathscr{L}_{\left[t,\xi_{\text{X}}-\mathcal{X}\right]}s.$ (214)
Note that the last two equations Eqs. (213) and (214) means that the perturbation $\left(0,0,\left[t,\xi_{\text{X}}-\mathcal{X}\right]^{a}\right)$ is trivial perturbation with trivial displacements $\left[t,\xi_{\text{X}}-\mathcal{X}\right]^{a}$. So equivalently, a linearized solution $\delta\phi$ is physically stationary if and only if there exists a smooth vector field $X^{a}$, which is an asymptotic symmetry near infinity, such that
$\mathscr{L}_{t}\delta\phi=$ $\left(-\mathscr{L}_{\left[t,\mathcal{X}\right]}g_{ab},-\mathscr{L}_{\left[t,\mathcal{X}\right]}A_{a},\zeta_{\text{X}}^{a}=\left[t,\mathcal{X}\right]^{a}\right)$
@@
Unfortunately, we will find that $\mathcal{E}$ is not degenerate on all physically stationary solutions. In fact, since $\mathcal{E}\left(\delta^{\prime}\phi,\delta\phi\right)=W\left(\delta^{\prime}\phi,\mathscr{L}_{t}\delta\phi\right)$, it follows that $\delta\phi$ is a degeneracy of $\mathcal{E}$ if and only if $\mathscr{L}_{t}\delta\phi$ is a degeneracy of $W$. As discussed at the end of Sec. V.3, when restricted to $\mathcal{C}$, $W$ is degenerate precisely on the gauge transformations $\mathscr{L}_{\mathcal{X}}\phi$ with $\mathcal{X}^{a}$ smooth and going to zero at infinity , so that $\delta\phi$ in the domain of $\mathcal{E}$ is a degeneracy if and only if
$\mathscr{L}_{t}\delta\phi=\left(\mathscr{L}_{\mathcal{X}}g_{ab},\mathscr{L}_{\mathcal{X}}A_{a},\xi_{\text{X}}^{a}=-\mathcal{X}^{a}\right).$ (217)
@@
$\mathcal{E}\left(\delta\phi,\delta_{\text{ps}}\phi\right)=$ $W\left(\delta\phi,\mathscr{L}_{t}\delta_{\text{ps}}\phi\right)$
$=$ $-W\left[\delta\phi,\mathscr{L}_{\left[t,\mathcal{X}\right]}\phi\right]$
@@
For a general asymptotic symmetry $\mathcal{X}^{a}$, the commutator $\left[t,\mathcal{X}\right]^{a}$ is, at most, an asymptotic translation (as occurs when $\mathcal{X}^{a}$ is an asymptotic boost). Therefore, in order the first term of right side vanishes, we need to restrict $\delta\phi$ such that $\delta H_{\left[t,\mathcal{X}\right]}=0$, where $\delta H_{\left[t,\mathcal{X}\right]}$ is the ADM linear momentum (see Eq. (46)). This in fact does not impose a physical restriction on the perturbations as we can achieve this by addition of the action of an infinitesimal Lorentz boost on the background solution. On the other hands, to make the second term vanishes, we need to restrict $\delta\phi$ such that
@@
for all trivial perturbations. As a consequence, let us define a subspace $\mathcal{V}\subset\mathcal{C}$ of the linearized solutions composed of perturbations that have $\delta H_{\left[t,\mathcal{X}\right]}=0$ and are symplectically orthogonal to all trivial perturbations, in other words, $\mathcal{V}$ is the symplectic complement of the subspace $\mathcal{W}$ spanned by the perturbations of the form $\mathscr{L}_{t}\delta_{\text{ps}}\phi$. Note that the double symplectic complement of $\mathcal{W}$ is itself according to Eq. (181), thus the symplectic complement of $\mathcal{V}$ in itself is $\mathcal{V}\cap\mathcal{W}$, which implies that the degeneracy of the canonical energy $\mathcal{E}$ is given precisely by the physically stationary solutions when restricted onto $\mathcal{V}$.
@@
Let us consider the restriction condition given in Eq. (219). For non-axisymmetric perturbations, this restricted condition is in fact same as the condition to make Eq. (201) vanish. If the superfluid neutrons and superconducting protons in our star are irrotational, i.e., their vorticity vanish $\bm{w}^{\Upsilon}=0$, and so that their total momentum covector should be of the form
@@
As shown in [32], if we further focus onto the background in which $\nabla_{a}s\neq 0$, then the condition Eq. (219) does not lead to a real physical restriction to $\mathcal{V}$ for non-axisymmetric perturbations, since it can be achieved in the suitable background solution by adding a trivial perturbation. However, if there are vortex in our superconducting-superfluid star and so that the irrotation is violated, then whether the condition Eq. (219) will impose a physical restriction to $\mathcal{V}$ is still unknown to our knowledge.
@@
$\Longleftrightarrow$ $\bm{\epsilon}^{abcd}\nabla_{a}\beta_{\text{X}}\bm{\epsilon}_{ebcd}n_{\text{X}}^{e}=\bm{\epsilon}^{abcd}\nabla_{a}\Omega_{\text{X}}\delta\left(\bm{\epsilon}_{ebcd}n_{\text{X}}^{e}\right),$
@@
So we still have the perturbed conservation of angular momentum, i.e., $\delta\left(d\bm{J}^{\text{X}}\right)=d\left(\delta\bm{J}^{\text{X}}\right)=0$. Replacing $\bm{N}^{\text{X}}$ by $\bm{J}^{\text{X}}$, then an identical calculation, besides the sixth equality in Eq. (229) replaced by
@@
where the second equality follows from the symmetry of $\mathcal{E}$, and the last equality follows from $\mathscr{L}_{t}\delta\phi$ satisfies $\Delta_{\text{X}}\bm{J}^{\text{X}}=0$. So $\mathscr{L}_{t}^{2}\delta\phi$ is symplectically orthogonal to all axisymmetric trivial perturbations. Furthermore, for a smooth vector field $\mathcal{X}^{a}$ which is an asymptotic symmetry near infinity, since
$W\left(\mathscr{L}_{\left[t,\mathcal{X}\right]}\phi,\mathscr{L}_{t}^{2}\delta\phi\right)=$ $\mathcal{E}\left(\mathscr{L}_{\left[t,\mathcal{X}\right]}\phi,\mathscr{L}_{t}\delta\phi\right)$
$=$ $\mathcal{E}\left(\mathscr{L}_{t}\delta\phi,\mathscr{L}_{\left[t,\mathcal{X}\right]}\phi\right)$
$=$ $W\left(\mathscr{L}_{t}\delta\phi,\mathscr{L}_{\left[t,\left[t,\mathcal{X}\right]\right]}\phi\right)$
@@
where we have used the fact that $\mathcal{Y}^{a}=\left[t,\left[t,\mathcal{X}\right]\right]^{a}$ vanishes at infinity for any asymptotic symmetry generator $\mathcal{X}^{a}$. Thus, we have shown that $\mathscr{L}_{t}^{2}\delta\phi\in\mathcal{V}$. ∎
@@ VII Thermodynamic stability
As shown in [19], there can be a non-axisymmetric perturbation which is made to have $\sum_{\text{X}}\Omega_{\text{X}}\delta^{2}J^{\text{X}}>\delta^{2}M$, whence $\mathcal{E}^{\prime}<0$, in other words, all rotating stars are thermodynamically unstable with respect to the non-axisymmetric perturbations. So we will solely consider the axisymmetric perturbations within the Lagrangian displacement framework such that $\delta J^{\text{X}}=0$ below, and we finally show that in the axisymmetric case, the canonical energy $\mathcal{E}\left(\delta\phi,\delta\phi\right)$ coincides with $\mathcal{E}^{\prime}$.
@@
$=$ $\int_{\Sigma}\left\{\delta\bm{\omega}^{\left(EM\right)}\left[\delta\phi,\left(\mathscr{L}_{t}g_{ab},\mathscr{L}_{t}A_{a}\right)\right]+\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right\}+\int_{\Sigma}\left[\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)-\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right]$
$=$ $\int_{\Sigma}\delta\bm{\omega}\left(\delta\phi,\mathscr{L}_{t}\phi\right)+\int_{\Sigma}\left[\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)-\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right]$
$=$ $\int_{\Sigma}\delta\left\{\iota_{t}\bm{E}\cdot\delta\phi+\delta\bm{C}_{t}+d\left[\delta\bm{Q}_{t}-\iota_{t}\bm{\theta}\left(\phi;\delta\phi\right)\right]\right\}+\int_{\Sigma}\left[\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)-\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)\right]$
@@
$\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)=\sum_{\text{X}}\left(\mathscr{L}_{t}\xi_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}-\xi_{\text{X}}^{a}\mathscr{L}_{t}\delta\bm{P}_{a}^{\text{X}}-\left[\xi_{\text{X}},\mathscr{L}_{t}\xi_{\text{X}}\right]^{a}\bm{P}_{a}^{\text{X}}\right),$ (245)
@@
$\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\delta\phi\right)-\delta\bm{\omega}^{\left(\text{M}\right)}\left(\delta\phi,\mathscr{L}_{t}\phi\right)=\sum_{\text{X}}\left(t^{a}\delta^{2}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{t}\xi_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}-\left[\xi_{\text{X}},\mathscr{L}_{t}\xi_{\text{X}}\right]^{a}\bm{P}_{a}^{\text{X}}\right).$ (247)
@@
$=$ $\sum_{\text{X}}\left[|v_{\text{X}}|u_{\text{X}}^{a}\delta^{2}\bm{P}_{a}^{\text{X}}-\Omega_{\text{X}}\varphi^{a}\delta^{2}\bm{P}_{a}^{\text{X}}-2\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\delta\bm{P}_{a}^{\text{X}}+\mathscr{L}_{\xi_{\text{X}}}\mathscr{L}_{\xi_{\text{X}}}\left(|v_{\text{X}}|u_{\text{X}}^{a}\right)\bm{P}_{a}^{\text{X}}\right].$ (250)
@@
$=$ $-\delta\left(\frac{1}{2}u_{\text{X}}^{a}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-q_{\text{X}b}^{a}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\bm{P}_{a}^{\text{X}}-\left(u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-2u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$
$=$ $-\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+\mathscr{L}_{\xi_{\text{X}}}\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+\left(u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-2u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$
$=$ $\mathscr{L}_{\xi_{\text{X}}}\left(\frac{1}{2}u_{\text{X}}^{a}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-q_{\text{X}b}^{a}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\bm{P}_{a}^{\text{X}}+\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\delta u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$
$=$ $\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}-\mathscr{L}_{\xi_{\text{X}}}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}-\left(\frac{1}{2}u_{\text{X}}^{b}u_{\text{X}}^{c}\delta g_{bc}-u_{\text{X}b}\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{b}\right)\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\bm{P}_{a}^{\text{X}}+2\mathscr{L}_{\xi_{\text{X}}}u_{\text{X}}^{a}\delta\bm{P}_{a}^{\text{X}}$